During the first part of a trip, a canoeist travels 90 miles at a certain speed. The canoeist travels 16 miles on the second part or the trip at a speed 5 mph slower. The total time for the trip is 5 hours. What is the speed on each part of the trip?

Answer by scott8148(6628)

let x="speed on 2nd part", so x+5="speed on 1st part" __ t=d/r

90/(x+5) + 16/x = 5 __ multiplying by x(x+5) __ 90x+16x+80=5x^2+25x

5x^2-81x-80=0 __ use quadratic formula to find x



ANSWER BELOW



x=17.13 (approx)